# Generating Directed Graphs with Dual Attention and Asymmetric Encoding

### v1

@@ 3.3 Graph transformer with dual attention
Graph Transformers dwivedi2021graphtransformer ; rampavsek2022recipe are effective at modeling node interactions while leveraging the adjacency matrix to encode structural information. However, in directed graphs, this matrix only captures source-to-target flow, overlooking potential reverse influences that can be equally important. Generating directed graphs thus requires models that (i) capture bidirectional information flow to learn how source and target nodes influence one another, and (ii) integrate structural signals across nodes, edges, and the global graph to preserve relational dependencies and ensure coherent outputs.
To address this, our model jointly processes stacked node features $\bm{X}$, edge features $\bm{E}$, and global features $\bm{y}$ through a Transformer composed of $L$ layers of a dual attention block (see Figure 2). This dual attention mechanism introduces two key innovations: (i) an attention-based aggregation scheme that jointly captures directional information, and (ii) modulation layers to incorporate information at different levels of granularity across nodes, edges, and the global graph structure.
@@
To model directional dependencies more effectively, we use a cross-attention mechanism between source-to-target edge features $\bm{E}_{\text{ST}}$ and their reversed counterparts $\bm{E}_{\text{TS}}$, which enables the model to reason bidirectionally across edges. By explicitly attending to both directions, our architecture better captures the complex, reciprocal relationships inherent in directed graphs, improving the expressiveness and accuracy of graph generation, which builds on ideas from wang2024directed . Concretely, we compute two directional attention maps between role-specific node projections:
@@
where $\bm{Q}_{\text{S}}$, $\bm{Q}_{\text{T}}$, $\bm{K}_{\text{S}}$, and $\bm{K}_{\text{T}}$ are the role-specific projections for source and target nodes, respectively, and $d_{q}$ is the query feature dimension. These attention weights are modulated using edge features through a FiLM layer (see Eq. 10), producing the updated attention maps $\bm{Y^{\prime}}_{\text{ST}}$ and $\bm{Y^{\prime}}_{\text{TS}}$. We use $\bm{Y^{\prime}}_{\text{ST}}$ to update the edge features, effectively combining local edge attributes with global graph information.
To consolidate directional information, we introduce an attention aggregation mechanism. Instead of treating the attentions from the two directions independently, we concatenate the modulated attention maps and apply a single softmax operation to obtain unified attention weights:
@@
where $h$ is the hidden dimension and the concatenation operation is performed along the node dimensions. These unified weights are then used to aggregate the value vectors $\bm{V}_{\text{S}}$ and $\bm{V}_{\text{T}}$ through weighted summation:
@@
This aggregation allows the model to assign asymmetric importance to the source-to-target and target-to-source directions by applying a joint softmax. The node features are then updated using a gated residual connection, which adaptively combines the original and updated features by learning a gate that controls how much of the new information should be integrated. Full technical details are provided in Appendix A
Figure 2: Network architecture of Directo. We stack $L$ dual attention layers that account for both source-to-target and target-to-source information via cross-attention mechanisms. $\bm{X}$, $\bm{E}$ and $\bm{y}$ denote the stacked input node, edge, and global features. $\bm{X}^{\prime}$, $\bm{E}^{\prime}$ and $\bm{y}^{\prime}$ are the output of the model, i.e., predicted clean node and edge distribution, and graph feature. FiLM perez2018film and PNA corso2020principal pooling layers are incorporated to enable flexible modulation between node, edge, and graph-level features.
@@
A graph denoising model should predict clean node and edge types from noisy inputs, which requires integrating both local interactions and global structure from node, edge, and graph-level features. Building on standard graph generation architectures vignac2023digress ; siraudin2024cometh ; qin2025defog , we incorporate Feature-wise Linear Modulation (FiLM) perez2018film and Principal Neighbourhood Aggregation (PNA) corso2020principal layers to enhance multi-scale feature fusion.
FiLM adaptively modulates edge features using attention signals: given the edge feature matrix $\bm{E}$, the attention matrix $\bm{E}_{\text{attn}}$, and the learnable weights $\bm{W}_{\text{FiLM}}^{1}$, $\bm{W}_{\text{FiLM}}^{2}$, it computes
@@
where $\odot$ denotes element-wise multiplication. These layers are also used to integrate global graph-level signals into the construction of edge representations.
To complement this, PNA layers aggregates multi-scale neighborhood information via pooling operations: given node features $\bm{X}\in\mathbb{R}^{n\times h}$ and a learnable weight matrix $\bm{W}_{\text{PNA}}$, PNA computes
@@
with concatenation performed along the feature dimension. We use PNA layers to update global graph representations based on node-level information at each attention pass.
Together, these components enable expressive and scalable integration of local and global features, allowing the model to capture higher-order structure critical for accurate graph denoising.
@@ 4 Benchmarking Directed Graph Generation
Despite growing interest in directed graphs sun2024towards , benchmarking remains underdeveloped bechlerspeicher2025positiongraphlearninglose . In particular, existing approaches to benchmarking digraph generation typically rely on performance in downstream tasks, failing to directly evaluate generative performance. With the objective of establishing standard datasets and metrics, we propose here a comprehensive benchmark that covers synthetic and real-world digraphs, along with tailored metrics for the digraph setting that directly assess key generative qualities, including sample validity, diversity, and distributional alignment.
@@ 4.1 Datasets
To enable systematic evaluation, we introduce a suite of synthetic directed datasets alongside two real-world use cases: (i) neural architecture search, where DAGs model computational pipelines, and (ii) scene graphs, where directed edges encode semantic relations in visual scenes.
@@
We sample digraphs from four distributions: Erdős–Rényi (binomial model) both general and DAGs variants; Price’s model, a directed analogue of the Barabási–Albert model that produces DAGs; and a directed version of the stochastic block model (SBM) dataset proposed in martinkus2022SPECTRE . Each dataset comprises 200 graphs, split into 128 training, 32 validation, and 40 test graphs.
@@
A relevant application of DAGs is Neural Architecture Search, which aims for an automated design of efficient Neural Networks elsken2019neural . For this task, we use the TPU Tiles dataset google2023tpugraphs as a benchmark. It consists of computational graphs extracted from the Tensor Processing Unit workloads. Following li2024layerdag , we split the data into 5,040 training, 630 validation, and 631 test graphs.
@@
Scene graphs structurally represent the semantics of visual scenes, supporting tasks such as image retrieval, captioning, and visual question answering in computer vision and NLP. Our focus is on generating new scene graphs, valuable for tasks such as data augmentation, scenario simulation, or enhancing model robustness, complementing recent work on scene-to-image generation johnson2018image ; yang2022diffusion . To accomplish this, we use the Visual Genome dataset krishna2017visualgenome , a widely adopted benchmark for scene graph-related tasks, providing richly annotated images with object labels, attributes, and pairwise relations. We extract a subset of these graphs, allocating 203 for training, 51 for validation, and 63 for testing. Each graph contains nodes representing objects, relationships, or attributes, each of them spanning several classes. Object nodes represent entities in the scene, relationship nodes mediate directed interactions between pairs of object nodes, and attribute nodes describe properties of individual object nodes.
Further details and statistics for all datasets are available in Appendix E.
@@ 4.2 Metrics
To evaluate our generations, we build upon the evaluation metrics from the undirected martinkus2022SPECTRE ; bergmeister2023efficient and directed law2024directedheatkernels graph generation literature, further extending them to the digraph generative setting.
@@
For synthetic datasets, assessing validity comprises statistical tests to check whether the sampled graphs follow the corresponding generative distribution and, in the case of DAG datasets, ensuring acyclicity. For the TPU Tiles dataset, we measure acyclicity, and, for the Visual Genome dataset, we verify the validation of the typed structural constraints (e.g., edges from relationship nodes can point only to object nodes).
@@
To measure generative diversity while avoiding trivial replication of the training data, we compute the uniqueness (fraction of non-isomorphic generated graphs) and novelty (fraction of generated graphs not in the training set). The V.U.N. ratio reports the proportion of samples that are simultaneously valid, unique, and novel.
Distribution alignment
@@
To evaluate proximity to the original distribution, we compute the Maximum Mean Discrepancy (MMD) between generated and test graphs across several graph statistics: incoming and outgoing degree distributions, directed clustering coefficients, spectral features, and wavelet-based features. In particular, spectral features are derived from the directed Laplacian chung2005laplacians :
$\bm{L}=\bm{I}-\frac{1}{2}\left(\bm{\Phi}^{\frac{1}{2}}\bm{P}\bm{\Phi}^{-\frac{1}{2}}+\bm{\Phi}^{-\frac{1}{2}}\bm{P}\bm{\Phi}^{\frac{1}{2}}\right),$ (12)
where $\bm{P}$ is the random walk transition matrix and $\bm{\Phi}$ is the diagonal matrix of the Perron vector of $\bm{P}$. This choice simplifies computations compared to alternatives such as the Magnetic Laplacian. We report normalized results as the ratio of statistics from generated samples to those from the training data, allowing for direct comparison. Finally, to summarize these ratios, we average them.
@@
Among all metrics, we consider the aggregated indicators, V.U.N. and the average graph statistics ratio, as the most informative. Additional details on the evaluation metrics are provided in Appendix F.
Erdős-Renyi Directed Acyclic Graph (ER-DAG)
Training set 0.0113 0.0103 0.0355 0.0038 0.0024 1.0 99.2 100 0 0
MLE 0.0083 $\pm$ 0.0004 0.0089 $\pm$ 0.0003 0.1318 $\pm$ 0.0077 0.0823 $\pm$ 0.0013 0.1162 $\pm$ 0.0018 15.1 $\pm$ 0.2 0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0 $\pm$ 0.0
D-VAE 0.6158 $\pm$ 0.0160 0.6246 $\pm$ 0.0103 1.0509 $\pm$ 0.00304 0.6160 $\pm$ 0.0315 0.5432 $\pm$ 0.0389 106.6 $\pm$ 5.4 0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0 $\pm$ 0.0
DiGress 0.0138 $\pm$ 0.0035 0.0143 $\pm$ 0.0050 0.1074 $\pm$ 0.0090 0.0073 $\pm$ 0.0017 0.0042 $\pm$ 0.0011 1.9 $\pm$ 0.3 34 $\pm$ 4.1 100 $\pm$ 0.0 100 $\pm$ 0.0 34 $\pm$ 4.1
DeFoG 0.0109 $\pm$ 0.0019 0.0108 $\pm$ 0.0017 0.0540 $\pm$ 0.0112 0.0061 $\pm$ 0.0008 0.0030 $\pm$ 0.0005 1.3 $\pm$ 0.1 67.5 $\pm$ 1.6 100 $\pm$ 0.0 100 $\pm$ 0.0 67.5 $\pm$ 1.6
Directo-DD RRWP 0.0142 $\pm$ 0.0040 0.0129 $\pm$ 0.0034 0.0408 $\pm$ 0.0053 0.0055 $\pm$ 0.0010 0.0048 $\pm$ 0.0014 1.4 $\pm$ 0.3 79 $\pm$ 3.7 100 $\pm$ 0.0 100 $\pm$ 0.0 79 $\pm$ 3.7
Directo-DD MagLap 0.0145 $\pm$ 0.0040 0.0134 $\pm$ 0.0033 0.0582 $\pm$ 0.0083 0.0063 $\pm$ 0.0015 0.0034 $\pm$ 0.0011 1.5 $\pm$ 0.2 85 $\pm$ 9.2 100 $\pm$ 0.0 100 $\pm$ 0.0 85 $\pm$ 9.2
Directo RRWP 0.0107 $\pm$ 0.0019 0.0104 $\pm$ 0.0012 0.1209 $\pm$ 0.0194 0.0054 $\pm$ 0.0005 0.0043 $\pm$ 0.0008 1.7 $\pm$ 0.1 94 $\pm$ 1.0 100 $\pm$ 0.0 100 $\pm$ 0.0 94 $\pm$ 1.0
Directo MagLap 0.0117 $\pm$ 0.0014 0.0110 $\pm$ 0.0015 0.0711 $\pm$ 0.0120 0.0055 $\pm$ 0.0016 0.0026 $\pm$ 0.0004 1.3 $\pm$ 0.2 92 $\pm$ 3.7 100 $\pm$ 0.0 100 $\pm$ 0.0 92 $\pm$ 3.7
Stochastic Bottleneck Model (SBM)
Training set 0.0031 0.0031 0.0274 0.0027 0.0011 1.0 97.7 100 0 0
MLE 0.0020 $\pm$ 0.0002 0.0020 $\pm$ 0.0002 0.1973 $\pm$ 0.0162 0.0087 $\pm$ 0.0003 0.0508 $\pm$ 0.0009 11.6 $\pm$ 0.2 0 $\pm$ 0.0 100 $\pm$ 0.0 100 $\pm$ 0.0 0 $\pm$ 0.0
DeFoG 0.0036 $\pm$ 0.0009 0.0034 $\pm$ 0.0008 0.1276 $\pm$ 0.0065 0.0345 $\pm$ 0.0028 0.0961 $\pm$ 0.0071 21.4 $\pm$ 1.6 13.5 $\pm$ 5.4 100 $\pm$ 0.0 100 $\pm$ 0.0 13.5 $\pm$ 5.4
Directo RRWP 0.0038 $\pm$ 0.0011 0.0037 $\pm$ 0.0012 0.0492 $\pm$ 0.0041 0.0035 $\pm$ 0.0007 0.0015 $\pm$ 0.0004 1.4 $\pm$ 0.2 87 $\pm$ 5.1 100 $\pm$ 0.0 100 $\pm$ 0.0 87 $\pm$ 5.1
@@ 5 Experiments
In this section, we first evaluate the flexibility of our method to generate directed graphs using our newly proposed benchmark, covering two of the synthetic and both real-world datasets. We then analyze the impact of our architectural improvements on generative performance.
@@
In the general directed setting, we compare Directo to a Maximum Likelihood Estimation (MLE) baseline over node count, node types, and edge types. The implementation details are provided in Appendix G.1. For DAG settings, we also include the only two publicly available baselines: D-VAE zhang2019dvae and LayerDAG li2024layerdag .
Notably, the architectural improvements introduced in Secs. 3.3 and 3.2 are agnostic to the underlying graph iterative refinement framework. To demonstrate their versatility and enable for wider experimentation, we extend Directo to a discrete diffusion vignac2023digress backbone (Directo-DD). Details of the adopted discrete diffusion formulation are provided in Sec. I.3. We also include as baselines both DeFoG qin2025defog and DiGress vignac2023digress , which are discrete flow matching and discrete diffusion methods designed for the undirected setting, respectively. To make them applicable to the directed setting, we adapt both models by removing their edge symmetrization operations.
For each experiment, we highlight the best result and second-best method. Further details on the experimental setup can be found in Appendix G.
5.1 Directed Graph Generation Performance

### v2

@@ 3.3 Graph transformer with dual attention
Graph Transformers (Dwivedi & Bresson, 2021; Rampášek et al., 2022) effectively model node interactions by leveraging the adjacency matrix to encode structural information. However, in directed graphs, it is crucial for the model to distinguish the different semantics carried by incoming and outgoing edges, which makes the standard graph transformer insufficiently expressive. Generating digraphs thus requires models that (i) capture bidirectional information flow to learn how source and target nodes influence each other distinctly, and (ii) integrate structural signals across nodes, edges, and the global graph to preserve relational dependencies and ensure coherent outputs.
To address this, our model jointly processes stacked node features $\bm{X}$, edge features $\bm{E}$, and global features $\bm{y}$ through a graph transformer composed of $L$ layers of a dual attention block (see Figure 2). This dual attention mechanism introduces two key components: (i) an attention-based aggregation scheme that captures directional information, and (ii) modulation layers to incorporate information at different levels of granularity across nodes, edges, and the global graph structure.
@@
To model directional dependencies more effectively, we use a cross-attention mechanism between source-to-target (outgoing) edge features $\bm{E}_{\text{ST}}$ and their reversed target-to-source (incoming) counterparts $\bm{E}_{\text{TS}}$. By explicitly attending to both directions, our architecture, which builds on ideas from Wang et al. (2024), processes information bidirectionally and better captures directed relationships, improving expressiveness and accuracy. Concretely, we compute two directional attention maps between role-specific node projections:
@@
where $\bm{Q}_{\text{S}}$, $\bm{Q}_{\text{T}}$, $\bm{K}_{\text{S}}$, and $\bm{K}_{\text{T}}$ are the role-specific projections for nodes that act as the source or the target of an edge, respectively, and $d_{q}$ is the query feature dimension. These attention weights are modulated through a FiLM layer (see Equation 8) using edge features to enable effective reasoning, producing the updated attention maps $\bm{Y^{\prime}}_{\text{ST}}$ and $\bm{Y^{\prime}}_{\text{TS}}$.
Figure 2: Network architecture of Directo. We stack $L$ dual attention layers that account for both source-to-target and target-to-source information via cross-attention mechanisms. $\bm{X}$, $\bm{E}$ and $\bm{y}$ denote the stacked input node, edge, and global features. $\bm{X}^{\prime}$, $\bm{E}^{\prime}$ and $\bm{y}^{\prime}$ are the output of the model, i.e., predicted clean node and edge distribution, and graph feature. FiLM (Perez et al., 2018) and PNA pooling layers (Corso et al., 2020) are incorporated to enable flexible modulation between node, edge, and graph-level features. Full technical details are provided in Appendix A.
To further consolidate directional information, we introduce an attention aggregation mechanism. Instead of treating the attentions from the two directions independently, we concatenate the modulated attention maps and apply a single softmax operation to obtain unified attention weights:
@@
where $h$ is the hidden dimension, $\bm{V}_{\text{S}},\bm{V}_{\text{T}}$ are the value vectors for source and target nodes, and the concatenation operation is performed along the node dimensions. These unified weights are then used to compute a convex (i.e., softmax-weighted) linear combination of the value vectors:
@@
This aggregation allows the model to assign asymmetric importance to the outgoing and the incoming directions. The node features are then updated using a gated residual connection, which adaptively combines the original and updated features by learning a gate that controls how much of the new information should be integrated in the update.
@@
A graph denoising model should predict clean node and edge types from noisy inputs, which requires integrating both local interactions and global structure from node, edge, and graph-level features. Building on standard graph generation architectures (Vignac et al., 2023a; Siraudin et al., 2024; Qin et al., 2025a), we incorporate Feature-wise Linear Modulation (FiLM) (Perez et al., 2018) and Principal Neighbourhood Aggregation (PNA) (Corso et al., 2020) layers to enhance multi-scale feature fusion. FiLM adaptively modulates edge features using attention signals: given the edge feature matrix $\bm{E}$, the attention matrix $\bm{E}_{\text{attn}}$, and the learnable weights $\bm{W}_{\text{FiLM}}^{1}$, $\bm{W}_{\text{FiLM}}^{2}$, it computes
@@
where $\odot$ denotes element-wise multiplication. These layers are also used to integrate global graph-level signals $\bm{y}$ into the construction of edge features.
To complement this, PNA layers aggregates multi-scale neighborhood information via pooling operations: given node (or edge) features $\bm{X}\in\mathbb{R}^{n\times h}$ and a learnable weight matrix $\bm{W}_{\text{PNA}}$:
@@
with concatenation performed along the feature dimension. We use PNA layers to update global graph signals based on local node- and edge-level information at each attention pass. Together, these components enable expressive and scalable integration of local and global features, allowing the model to capture higher-order structure which is critical for accurate graph denoising.
@@ 4 Benchmarking Directed Graph Generation
Despite growing interest in digraphs (Sun et al., 2024), benchmarking in this area remains underdeveloped (Bechler-Speicher et al., 2025). Existing approaches typically rely on downstream task performance, failing to directly evaluate generative quality. With the objective of establishing standard datasets and metrics, we propose a comprehensive benchmark that covers synthetic and real-world digraphs, along with tailored metrics that directly assess key generative qualities, including sample validity, diversity, and distributional alignment.
@@ 4.1 Datasets
To enable systematic evaluation, we introduce a suite of synthetic directed datasets alongside two real-world use cases: (i) neural architecture search (NAS), where DAGs model computational pipelines, and (ii) scene graphs, where directed edges encode semantic relations in visual scenes. Further details and statistics for all datasets are available in Appendix E.
@@
We sample from the directed and DAG Erdős–Rényi (binomial) model; from Price’s model, a directed analogue of the Barabási–Albert model that produces DAGs; and from a directed version of the Stochastic Block Model (SBM) dataset proposed in Martinkus et al. (2022). Each dataset comprises 200 graphs, split into 128 training, 32 validation, and 40 test graphs.
@@
A relevant application of DAGs is Neural Architecture Search, which aims to generate novel instances to achieve an automated design of efficient Neural Networks (Elsken et al., 2019). To support this task, the TPU Tiles dataset (Phothilimthana et al., 2023) contains computational graphs extracted from Tensor Processing Unit workloads. We split into 5,040 training, 630 validation, and 631 test graphs following Li et al. (2025).
@@
Scene graphs represent the semantics of visual scenes, supporting tasks such as image retrieval, captioning, or visual question answering. Generating such graphs is useful for data augmentation, scenario simulation, or robustness, complementing recent scene-to-image generation work (Johnson et al., 2018; Yang et al., 2022). We use the Visual Genome dataset (Krishna et al., 2017), a widely adopted benchmark of graphs that are not acyclic but instead represent general directed structures, with nodes corresponding to objects (entities), relationships (directed interactions), or attributes (object properties). We extract 203 training, 51 validation, and 63 test graphs.
@@ 4.2 Metrics
To evaluate digraph generations, we build upon the evaluation metrics from the undirected graph generation literature  (Martinkus et al., 2022; Bergmeister et al., 2024; Thompson et al., 2022) and further extend the ones already present in the directed setting (Law et al., 2025).
@@
For synthetic datasets, validity is assessed via statistical tests of adherence to the generative distribution and, for DAG datasets, by ensuring acyclicity. On the TPU Tiles dataset, we measure acyclicity, while for Visual Genome we check typed structural constraints (e.g., edges go from objects to attributes and relationships, and from relationships to objects). To measure generative diversity while avoiding memorization, we compute uniqueness (fraction of non-isomorphic generated graphs) and novelty (fraction of generated graphs not in the training set). The V.U.N. ratio reports the proportion of samples that are simultaneously valid, unique, and novel.
Structural distributional alignment
@@
To evaluate proximity to the original distribution, we compute the Maximum Mean Discrepancy (MMD) (Martinkus et al., 2022) between generated and test graphs across several statistics: outgoing and incoming degree distributions, directed clustering coefficients, spectral, and wavelet-based features. Spectral features are derived from the directed Laplacian (Chung, 2005), simplifying computations compared to the Magnetic Laplacian. We normalize results as the ratio from the generated samples MMD to that of the training data and then average the ratios.
Joint node-edge distributional alignment
To better capture performance on attributed graphs, we extend evaluation beyond structural statistics to joint node–edge distributions. Specifically, we measure MMD over label histograms to assess coverage of categorical proportions. We also compute triplet-based precision and recall by matching semantic tuples (e.g., $(\text{object},\text{relation},\text{object})$) between generated and test graphs. Finally, we evaluate embedding-based distances (MMD with RBF kernel and FID) over GNN-derived representations that encode both labels and structure (Thompson et al., 2022). These metrics provide a comprehensive view of alignment with the reference distribution.
@@
Among all metrics, we consider V.U.N. and the average ratio as the most informative for structural distributional alignment, and the RBF MMD for the joint node-label distributional alignment in real-world datasets. Additional details on the evaluation metrics are provided in Appendix F.
@@ 5 Experiments
In this Section, we first evaluate the flexibility of our method to generate digraphs across synthetic and real-world datasets. Then, we analyze the role of the proposed architectural improvements on generative performance and analyze scalability and conditional generation.
@@
In the general directed setting, we compare Directo against performing Maximum Likelihood Estimation (MLE) baseline of the node count, node types, and edge types, and sampling from the resulting distribution to generate new graph instances. For DAG settings, we also include the only two publicly available baselines: D-VAE (Zhang et al., 2019) and LayerDAG (Li et al., 2025). Notably, the architectural improvements in Sections 3.2 and 3.3 are agnostic to the underlying iterative refinement framework. To demonstrate versatility and enable wider experimentation, we extend Directo to a discrete diffusion (Vignac et al., 2023a) backbone (Directo-DD), described in Section I.3. We also include DeFoG (Qin et al., 2025a) and DiGress (Vignac et al., 2023a), discrete flow matching and diffusion methods in the undirected setting, respectively, adapting both to the directed case by removing the edge symmetrization operations. For each experiment, we highlight the best result and second-best method, with further experimental details in Appendix G.
5.1 Generative Performance Evaluation
Table 1: Directed graph generation performance for different configurations of Directo. Results are the mean $\pm$ standard deviation across five sampling runs. We considered $Q=10$ for MagLap in the synthetic datasets and $Q=5$ in the real-world ones (due to the computational cost of this positional encoding). OOT indicates that the model could not be run within a reasonable timeframe.
ER-DAG SBM TPU Tiles Visual Genome
Model Ratio $\downarrow$ V.U.N. $\uparrow$ Ratio $\downarrow$ V.U.N. $\uparrow$ Ratio $\downarrow$ V.U.N. $\uparrow$ RBF $\downarrow$ Ratio $\downarrow$ V.U.N. $\uparrow$ RBF $\downarrow$
Training set 1.0 0.0 1.0 0.0 1.0 0.0 0.002 1.0 0.0 0.021
MLE 15.1 $\pm$ 0.2 0.0 $\pm$ 0.0 11.6 $\pm$ 0.2 0.0 $\pm$ 0.0 149.8 $\pm$ 0.7 24.7 $\pm$ 0.0 1.039 $\pm$ 0.033 17.0 $\pm$ 0.6 0.0 $\pm$ 0.0 0.618 $\pm$ 0.025
D-VAE 106.6 $\pm$ 5.4 0.0 $\pm$ 0.0 - - OOT OOT OOT - - -
LayerDAG 4.2 $\pm$ 3.2 21.5 $\pm$ 2.7 - - 413.6 $\pm$ 70.1 98.5 $\pm$ 3.0 1.021 $\pm$ 0.023 - - -
DiGress 1.9 $\pm$ 0.3 34.0 $\pm$ 4.1 3.9 $\pm$ 0.9 41.5 $\pm$ 5.1 57.5 $\pm$ 1.7 70.9 $\pm$ 3.4 0.097 $\pm$ 0.033 17.0 $\pm$ 0.6 0.3 $\pm$ 0.6 0.232 $\pm$ 0.028
DeFoG 1.6 $\pm$ 0.2 75.0 $\pm$ 2.2 4.3 $\pm$ 0.8 37.0 $\pm$ 6.6 63.7 $\pm$ 2.6 72.0 $\pm$ 2.4 0.059 $\pm$ 0.015 10.8 $\pm$ 0.7 50.8 $\pm$ 8.4 0.085 $\pm$ 0.023
Directo-DD RRWP 1.4 $\pm$ 0.3 79.0 $\pm$ 3.7 1.7 $\pm$ 0.4 81.5 $\pm$ 3.2 61.0 $\pm$ 2.9 76.8 $\pm$ 1.9 0.058 $\pm$ 0.023 15.3 $\pm$ 0.8 72.7 $\pm$ 3.9 0.039 $\pm$ 0.004
Directo-DD MagLap 1.5 $\pm$ 0.2 85.0 $\pm$ 9.2 1.5 $\pm$ 0.4 95.5 $\pm$ 3.7 64.3 $\pm$ 5.3 77.0 $\pm$ 7.0 0.079 $\pm$ 0.027 7.6 $\pm$ 0.7 61.9 $\pm$ 4.4 0.042 $\pm$ 0.006
Directo RRWP 1.7 $\pm$ 0.1 94.0 $\pm$ 1.0 1.8 $\pm$ 0.5 99.5 $\pm$ 1.0 75.4 $\pm$ 8.1 77.0 $\pm$ 2.9 0.044 $\pm$ 0.018 12.8 $\pm$ 0.6 83.8 $\pm$ 4.3 0.038 $\pm$ 0.005
Directo MagLap 1.3 $\pm$ 0.2 92.0 $\pm$ 3.7 2.0 $\pm$ 0.3 96.5 $\pm$ 2.5 44.0 $\pm$ 7.1 80.5 $\pm$ 4.6 0.042 $\pm$ 0.001 6.2 $\pm$ 0.5 67.0 $\pm$ 4.3 0.051 $\pm$ 0.012